The Linux XFree86 HOWTO

by Eric S. Raymond v4.0, 10 November 1996 This document describes how to obtain, install, and configure version 3.2 of the XFree86 version of the X Window System (X11R6) for Linux systems. It is a step-by-step guide to configuring XFree86 on your system.

Contents

1 Introduction 1 1.1 Other sources of information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 New versions of this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Feedback and Corrections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Hardware requirements 2

3 Installing XFree86 5

4 Configuring XFree86 8 4.1 Normal Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 Troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.3 Custom Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

5 Running XFree86 10 5.1 Terms of Use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.2 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

1 Introduction

The X Window System is a large and powerful (some might say excessively large and overly complex) graphics environment for UNIX systems. The original X Window System code was developed at MIT; commercial vendors have since made X the industry standard for UNIX platforms. Virtually every UNIX workstation in the world runs some variant of the X Window system. A freely redistributable port of the MIT X Window System version 11, release 6 (X11R6) for 80386/80486/Pentium UNIX systems has been developed by a team of programmers originally headed by David Wexelblat <dwex@XFree86.org>. The release, known as XFree86, is available for System V/386, 386BSD, and other x86 UNIX implementations, including Linux. It includes all of the required binaries, support files, libraries, and tools.

Complete information on XFree86 is available at the XFree86 web site, <http://www.XFree86.org>.

In this document, we’ll give a step-by-step description of how to install and configure XFree86 for Linux, but you will have to fill in some of the details yourself by reading the documentation released with XFree86 itself. (This documentation is discussed below.) However, using and customizing the X Window System is far beyond the scope of this document—for this purpose you should obtain one of the many good books on using the X Window System.

1.1 Other sources of information

If you have never heard of Linux before, there are several sources of basic information about the system. The best place to find these is at the Linux Documentation Project home page at <http://sunsite.unc.edu/mdw/linux.html>. You can find the latest, up-to-date version of this

document there, as <http://sunsite.unc.edu/mdw/HOWTO/XFree86-HOWTO.html>

1.2 New versions of this document

New versions of the Linux XFree86 HOWTO will be periodically posted to comp.os.linux.help and and news.answers <news:answers>. They will also be uploaded to various Linux WWW and FTP sites, including the LDP home page.

You can always view the latest version of this on the World Wide Web via the URL

<http://sunsite.unc.edu/mdw/HOWTO/XFree86-HOWTO.html>.

1.3 Feedback and Corrections

If you have questions or comments about this document, please feel free to mail Eric S. Raymond, at esr@thyrsus.com. I welcome any suggestions or criticisms. If you find a mistake with this document, please let me know so I can correct it in the next version. Thanks.

Please do not mail me questions about how to make your video card and monitor work with X. This HOWTO is intended to be a rapid, painless guide to normal installation using the new interactive configurator. If you run into problems, browse the Hitchiker’s Guide to XFree86 Video Timing,

<http://ww.ccil.org/ esr/xconfig/video-modes.html>. (This is the up-to-date HTML version of

XFree86’s ‘Videomodes.doc’ file.) That document tells everything I know about configuration troubleshooting. If it can’t help you, I can’t either.

2 Hardware requirements

As of XFree86 version 3.2, released November 1996, the following video chipsets are supported. The documentation included with your video adaptor should specify the chipset used. If you are in the market for a new video card, or are buying a new machine that comes with a video card, have the vendor find out exactly what the make, model, and chipset of the video card is. This may require the vendor to call technical support on your behalf; in general vendors will be happy to do this. Many PC hardware vendors will state that the video card is a “standard SVGA card” which “should work” on your system. Explain that your software (mention Linux and XFree86!) does not support all video chipsets and that you must have detailed information. You can also determine your videocard chipset by running the SuperProbe program included with the XFree86 distribution. This is covered in more detail below. The following standard SVGA chipsets are supported:

• Tseng ET3000, ET4000AX, ET4000/W32, ET6000

• Western Digital/Paradise PVGA1

• Western Digital WD90C00, WD90C10, WD90C11, WD90C24, WD90C30, WD90C31, WD90C33 • Genoa GVGA

• Trident TVGA8800CS, TVGA8900B, TVGA8900C, TVGA8900CL, TVGA9000, TVGA9000i, TVGA9100B, TVGA9200CX, TVGA9320, TVGA9400CX, TVGA9420, TGUI9420DGi, TGUI9430DGi, TGUI9440AGi, TGUI9660XGi, TGUI9680

• ATI 18800, 18800-1, 28800-2, 28800-4, 28800-5, 28800-6, 68800-3, 68800-6, 68800AX, 68800LX, 88800GX-C, 88800GX-D, 88800GX-E, 88800GX-F, 88800CX, 264CT, 264ET, 264VT, 264VT2, 264GT

• NCR 77C22, 77C22E, 77C22E+

• Cirrus Logic CLGD5420, CLGD5422, CLGD5424, CLGD5426, CLGD5428, CLGD5429, CLGD5430, CLGD5434, CLGD5436, CLGD5440, CLGD5446, CLGD5462, CLGD5464, CLGD6205, CLGD6215, CLGD6225, CLGD6235, CLGD6410, CLGD6412, CLGD6420, CLGD6440

• OAK OTI067, OTI077, OTI087

• Avance Logic ALG2101, ALG2228, ALG2301, ALG2302, ALG2308, ALG2401

• Chips & Technologies 65520, 65530, 65540, 65545, 65520, 65530, 65540, 65545, 65546, 65548, 65550, 65554

• MX MX68000, MX680010

• Video 7/Headland Technologies HT216-32

• SiS 86C201, 86C202, 86C205

• ARK Logic ARK1000PV, ARK1000VL, ARK2000PV, ARK2000MT

• RealTek RTG3106

• Alliance AP6422

• Matrox MGA2064W

• NVidia/SGS Thomson NV1, STG2000 The following SVGA chipsets with accelerated features are also supported:

• 8514/A (and true clones)

• ATI Mach8, Mach32, Mach64

• Cirrus CLGD5420, CLGD5422, CLGD5424, CLGD5426, CLGD5428, CLGD5429, CLGD5430, CLGD5434, CLGD5436, CLGD5440, CGLD5446, CLGD5462, CLGD5464.

• S3 86C911, 86C924, 86C801, 86C805, 86C805i, 86C928, 86C864, 86C964, 86C732, 86C764, 86C765, 86C868, 86C968, 86C325, 86C988

• Western Digital WD90C31, WD90C33, WD90C24A

• Weitek P9000

• IIT AGX-014, AGX-015, AGX-016

• IBM XGA-2

• Tseng ET4000/W32, ET4000/W32i, ET4000/W32p, ET6000

• Ark Logic ARK1000PV, ARK1000VL, ARK2000PV, ARK2000MT

• MGA2064W

Video cards using these chipsets are supported on all bus types, including VLB and PCI.

All of the above are supported in both 256 color and monochrome modes, with the exception of the Avance Logic, MX and Video 7 chipsets, which are only supported in 256 color mode. If your video card has enough DRAM installed, many of the above chipsets are supported in 16 and 32 bits-per-pixel mode (specifically, some Mach32, P9000, S3 and Cirrus boards). The usual configuration is 8 bits per pixel (that is, 256 colors).

The monochrome server also supports generic VGA cards, the Hercules monochrome card, the Hyundai HGC1280, Sigma LaserView, and Apollo monochrome cards. On the Compaq AVGA, only 64k of video memory is supported for the monochrome server, and the GVGA has not been tested with more than 64k.

This list will undoubtedly expand as time passes. The release notes for the current version of XFree86 should contain the complete list of supported video chipsets.

One problem faced by the XFree86 developers is that some video card manufacturers use non-standard mechanisms for determining clock frequencies used to drive the card. Some of these manufacturers either don’t release specifications describing how to program the card, or they require developers to sign a non-disclosure statement to obtain the information. This would obviously restrict the free distribution of the XFree86 software, something that the XFree86 development team is not willing to do. For a long time, this has been a problem with certain video cards manufactured by Diamond, but as of release 3.1 of XFree86, Diamond has started to work with the development team to release free drivers for these cards.

The suggested setup for XFree86 under Linux is a 486 machine with at least 8 megabytes of RAM, and a video card with a chipset listed above. For optimal performance, we suggest using an accelerated card, such as an S3-chipset card. You should check the documentation for XFree86 and verify that your particular card is supported before taking the plunge and purchasing expensive hardware. Benchmark ratings comparisons for various video cards under XFree86 are posted routinely to the USENET newsgroups comp.windows.x.i386unix <news:comp.windows.x.i386unix> and comp.os.linux.x

<news:comp.os.linux.x>.

As a side note, the personal Linux system of Matt Welsh (this FAQ’s originator) was a 486DX2-66, 20 megabytes of RAM, equipped with a VLB S3-864 chipset card with 2 megabytes of DRAM. He ran X benchmarks on this machine as well as on Sun Sparc IPX workstations. The Linux system was roughly 7 times faster than the Sparc IPX (for the curious, XFree86-3.1 under Linux, with this video card, runs at around 171,000 xstones; the Sparc IPX at around 24,000). In general, XFree86 on a Linux system with an accelerated SVGA card will give you much greater performance than that found on commercial UNIX workstations (which usually employ simple framebuffers for graphics).

Your machine will need at least 4 megabytes of physical RAM, and 16 megabytes of virtual RAM (for example, 8 megs physical and 8 megs swap). Remember that the more physical RAM that you have, the less that the system will swap to and from disk when memory is low. Because swapping is inherently slow (disks are very slow compared to memory), having 8 megabytes of RAM or more is necessary to run XFree86 comfortably. 16 is better. A system with 4 megabytes of physical RAM could run much (up to 10 times) more slowly than one with 8 megs or more.

3 Installing XFree86

It’s quite likely that you obtained XFree86 as part of a Linux distribution, in which case downloading the software separately is not necessary. In that case you can skip this section.

The Linux binary distribution of XFree86 can be found on a number of FTP sites. On the XFree86 site it’s

at <ftp://ftp.xfree86.org/pub/XFree86/current/binaries/Linux> (As of the time of this writing,

the current version is 3.2; newer versions are released periodically).

On sunsite.unc.edu, it will be found in the directory /pub/Linux/X11.

If you are downloading XFree86 directly, this table lists the files in the XFree86-3.2 distribution.

One of the following servers is required:

X328514.tgz Server for 8514-based boards.

X32AGX.tgz Server for AGX-based boards.

X32I128.tgz Server for the Number Nine Imagine 128.

X32Mach32.tgz Server for Mach32-based boards.

X32Mach64.tgz

Server for Mach64-based boards. X32Mach8.tgz

Server for Mach8-based boards. X32Mono.tgz

Server for monochrome video modes. X32P9K.tgz

Server for P9000-based boards. X32S3.tgz

Server for S3-based boards. X32S3V.tgz

Server for the S3 ViRGE and ViRGE/VX (considered beta) X32SVGA.tgz

Server for Super VGA-based boards. X32W32.tgz

Server for ET4000/W32-based boards.

If you don’t know which one to take, take the VGA16 server, X32VGA16.tgz. You want to download this one anyway, because you’ll need it to run the auto-configuration utility in the next step. All of the following files are required:

preinst.sh

Pre-installation script postinst.sh

Post-installation script X32bin.tgz

The rest of the X11R6 binaries. X32cfg.tgz

Config files for xdm, xinit and fs. X32doc.tgz

Documentation and manpages. X32fnts.tgz

75dpi, misc and PEX fonts X32lib.tgz

Shared X libraries and support files. X32set.tgz

XF86Setup utility X32VG16.tgz

Server for VGA/EGA-based boards.

The following files are optional:

X32f100.tgz

100dpi fonts X32fcyr.tgz

Cyrillic fonts X32fnon.tgz

Other fonts (Chinese, Japanese, Korean, Hebrew) X32fscl.tgz

Scalable fonts (Speedo and Type1) X32fsrv.tgz

Font server and config files X32prog.tgz

X header files, config files and compile-time libs X32lkit.tgz

X server LinkKit X32lk98.tgz

PC98 X server LinkKit X32nest.tgz

Nested X server X32vfb.tgz

Virtual framebuffer X server X32ps.tgz

PostScript version of the documentation X32html.tgz

HTML version of the documentation The XFree86 directory should contain release notes for the current version in RELNOTES. Consult those for installation details

All that is required to install XFree86 is to obtain the above files, create the directory /usr/X11R6 (as root), and unpack the files from /usr/X11R6 with a command such as:

     gzip -dc X32bin.tgz | tar xfB -

Remember that these tar files are packed relative to /usr/X11R6, so it’s important to unpack the files there.

You need to make sure that /usr/X11R6/bin is on your path. This can be done by editing your system default /etc/profile or /etc/csh.login (based on the shell that you, or other users on your system, use). Or you can simply add the directory to your personal path by modifying /etc/.bashrc or /etc/.cshrc, based on your shell.

You also need to make sure that /usr/X11R6/lib can be located by ld.so, the runtime linker. To do this, add the line

     /usr/X11R6/lib
to the file /etc/ld.so.conf, and run /sbin/ldconfig, as root.

Note: I found that, after installing all of the following parts:

X32Ma64.tar X32cfg.tar X32fnts.tar X32html.tar X32prog.tar X32VG16.tar X32doc.tar X32fscl.tar X32lib.tar X32set.tar X32bin.tar X32f100.tar X32fsrv.tar X32man.tar

that the Xpm support wasn’t included in any of them (this is a problem because the X server itself requires Xpm, at lease in my configuration; your mileage may vary). I haad to copy the Xpm libraries from my old X distribution. So the packaging isn’t perfect yet.

4 Configuring XFree86

4.1 Normal Configuration

Configuring XFree86 to use your mouse, keyboard, monitor, and video card correctly used to be something of a black art, requiring extensive hand-hacking of a complex configuration file. No more; the 3.2 release has made the process nearly trivial. All you do is fire up the program XF86Setup.

This program depends on the fact that all new PC hardware these days ships with EGA/VGA capable monitors. It invokes the SVGA16 server and uses it to bring up X in a lowest-common-denominator 640x480 mode. Then it runs an interactive program that walks you through a series of five configuration panels – mouse, keyboard, (video) card, monitor, and ‘other’ (miscellaneous server options). The whole process is quite painless.

One minor point to keep in mind is that, if you’re like most people using current PC, your keyboard is actually what XF86Setup calls ‘Generic 102-key PC (intl)’ rather than the default ‘Generic 101-key PC’. If you pick the default (101) the key cluster on the extreme right of your keyboard (numeric keypad and friends) may stop working.

If you’re not sure of your monitor type, you can try the listed ones in succession. Work your way from top down (upper choices involve lower dot-clock speeds and are less demanding on the hardware). Back off if you get hash or a seriously distorted picture. Minor distortions (picture slightly too large, slightly too small, or slightly off-center) are no problem; you’ll get a chance to correct those immediately by fine-tuning the mode.

And, when the program brings up xvidtune to allow you to tweak your video mode, don’t let the initial warning box make you nervous. Modern multisync monitors (unlike their fixed-frequency predecessors) are not easy to damage this way.

4.2 Troubleshooting

Occasionally, something will not be quite right when you initially fire up the X server. This is almost always caused by a problem in your configuration file. Usually, the monitor timing values are off, or the video card dot clocks set incorrectly. Minor problems can be fixed with xvidtune; a really garbled screen usually means you need to go back into XF86Setup and choose a less capable monitor type.

If your display seems to roll, or the edges are fuzzy, this is a clear indication that the monitor timing values or dot clocks are wrong. Also be sure that you are correctly specifying your video card chipset, as well as other options for the Device section of XF86Config. Be absolutely certain that you are using the right X server and that /usr/X11R6/bin/X is a symbolic link to this server.

If all else fails, try to start X “bare”; that is, use a command such as:

     X > /tmp/x.out 2>&1

You can then kill the X server (using the ctrl-alt-backspace key combination) and examine the contents of /tmp/x.out. The X server will report any warnings or errors—for example, if your video card doesn’t have a dot clock corresponding to a mode supported by your monitor.

Remember that you can use ctrl-alt-numeric + and ctrl-alt-numeric - to switch between the video modes listed on the Modes line of the Screen section of XF86Config. If the highest resolution mode doesn’t look right, try switching to lower resolutions. This will let you know, at least, that those parts of your X configuration are working correctly.

Also, check the vertical and horizontal size/hold knobs on your monitor. In many cases it is necessary to adjust these when starting up X. For example, if the display seems to be shifted slightly to one side, you can usually correct this using the monitor controls.

The USENET newsgroup comp.windows.x.i386unix is devoted to discussions about XFree86, as is comp.os.linux.x. It might be a good idea to watch that newsgroup for postings relating to your video configuration—you might run across someone with the same problems as your own.

4.3 Custom Configuration

You will need to hand-hack your X configuration to get optimal performance if your monitor can support 1600x1200 – the highest canned resolution XF86Setup supports is 1280x1024.

If you want to hand-hack your video configuration for this or any other reason, go see the Hitchiker’s Guide

to XFree86 Video Timing, <http://ww.ccil.org/ esr/xconfig/video-modes.html>. (This is the

up-to-date HTML version of XFree86’s ‘Videomodes.doc’ file.)

5 Running XFree86

With your XF86Config file configured, you’re ready to fire up the X server and give it a spin. First, be sure that /usr/X11R6/bin is on your path.

The command to start up XFree86 is

     startx

This is a front-end to xinit (in case you’re used to using xinit on other UNIX systems).

This command will start the X server and run the commands found in the file .xinitrc in your home directory. .xinitrc is just a shell script containing X clients to run. If this file does not exist, the system

default /usr/X11R6/lib/X11/xinit/xinitrc will be used.

A standard .xinitrc file looks like this:

     #!/bin/sh
     xterm -fn 7x13bold -geometry 80x32+10+50 &
     xterm -fn 9x15bold -geometry 80x34+30-10 &
     oclock -geometry 70x70-7+7 &
     xsetroot -solid midnightblue &
     exec twm

This script will start up two xterm clients, an oclock, and set the root window (background) color to midnightblue. It will then start up twm, the window manager. Note that twm is executed with the shell’s exec statement; this causes the xinit process to be replaced with twm. Once the twm process exits, the X server will shut down. You can cause twm to exit by using the root menus: depress mouse button 1 on the desktop background—this will display a pop up menu which will allow you to Exit Twm.

Be sure that the last command in .xinitrc is started with exec, and that it is not placed into the background (no ampersand on the end of the line). Otherwise the X server will shut down as soon as it has started the clients in the .xinitrc file.

Alternately, you can exit X by pressing ctrl-alt-backspace in combination. This will kill the X server directly, exiting the window system.

The above is a very, very simple desktop configuration. Many wonderful programs and configurations are available with a bit of work on your .xinitrc file. For example, the fvwm window manager will provide a virtual desktop, and you can customize colors, fonts, window sizes and positions, and so forth to your heart’s content.

If you are new to the X Window System environment, we strongly suggest picking up a book such as The X Window System: A User’s Guide. Using and configuring X is far too in-depth to cover here. See the man pages for xterm, oclock, and twm for clues on getting started.

5.1 Terms of Use

This document is copyright 1996 by Eric S. Raymond. You may use, disseminate, and reproduce it freely, provided you:

• Do not omit or alter this copyright notice.

• Do not omit or alter or omit the version number and date.

• Do not omit or alter the document’s pointer to the current WWW version.

• Clearly mark any consdensed, altered or versions as such.

These restrictions are intended to protect potential readers from stale or mangled versions. If you think you have a good case for an exception, ask me.

5.2 Acknowledgements

This document was originated by Matt Welsh in the dim and backward abysm of time. Thanks, Matt!